# Reframing the Relationship in Out-of-Distribution Detection

### v1

@@ 4 Experiments
MSP [24] 40.89 88.63 65.81 81.24 67.90 80.14 64.96 78.16 59.89 82.04
Energy [45] 21.59 95.99 34.28 93.15 36.64 91.82 51.18 88.09 35.92 92.26
ODIN [42] 30.22 94.65 54.04 87.17 55.06 85.54 51.67 87.85 47.75 88.80
ViM [71] 32.19 93.16 54.01 87.19 60.67 83.75 53.94 87.18 50.20 87.82
KNN [65] 29.17 94.52 35.62 92.67 39.61 91.02 64.35 85.67 42.19 90.97
NPOS [66] 16.58 96.19 43.77 90.44 45.27 89.44 46.12 88.80 37.93 91.22
CoOp [86] 43.38 91.26 38.53 91.95 46.68 89.09 50.64 87.83 44.81 90.03
LoCoOp [52] 38.49 92.49 33.27 93.67 39.23 91.07 49.25 89.13 40.17 91.53
@@
FashionMNIST [76] 0.00 100.00 0.00 100.00 0.00 100.00 13.61 93.60 3.40 98.40
STL10 [10] 0.00 100.00 2.56 99.01 0.00 100.00 0.00 100.00 0.64 99.75
OxfordIIIPet [55] 0.00 99.89 0.37 99.89 0.96 99.71 0.35 99.85 0.42 99.84
Food101 [6] 0.25 99.90 0.49 99.86 1.79 99.40 2.99 99.33 1.38 99.62
CUB-200 [69] 0.00 100.00 0.00 100.00 0.00 99.99 0.00 100.00 0.00 100.00
PlantVillage [30] 2.95 97.83 0.18 98.11 1.12 98.41 4.37 97.56 2.16 97.98
LFW [27] 2.89 99.14 2.24 99.52 8.04 98.25 18.88 95.34 8.01 98.06
Stanford-dogs [34] 0.11 99.92 0.16 99.91 0.58 99.73 0.67 99.72 0.38 99.82
FGVC-Aircraft [47] 0.00 99.99 0.67 99.87 1.02 99.69 0.00 99.99 0.42 99.89
Grocery Store [36] 0.03 99.89 0.15 99.97 0.69 99.82 0.79 99.76 0.42 99.86
CIFAR10 [37] 0.00 100.00 5.12 98.29 2.56 99.67 0.00 99.88 1.92 99.46
CIFAR100 [37] 12.80 97.15 17.92 95.66 15.36 96.19 4.53 98.42 12.65 96.86
Table 3: Comparison results on hard OOD detection tasks.
FPR95 / AUROC FPR95 / AUROC FPR95 / AUROC FPR95 / AUROC FPR95 / AUROC
MSP [24] 9.38 / 98.31 12.51 / 97.70 5.82 / 98.88 14.14 / 95.73 39.57 / 90.99
MCM [48] 5.00 / 98.71 12.91 / 98.09 3.70 / 99.09 13.16 / 96.32 5.87 / 98.36
CMA(Ours) 3.1 / 99.19 6.2 / 98.71 2.94 / 99.18 10.32 / 96.64 3.22 / 99.01
@@ 4.1 Setup
We conducted a comprehensive evaluation of the performance of our method across various dimensions and compared it with widely employed OOD detection algorithms. (1) We assessed our approach on the ImageNet-1k OOD benchmark. This benchmark utilizes the large-scale visual dataset ImageNet-1k [13] as the ID data and four OOD datasets (including subsets of iNaturalist [67], SUN [77], Places[85], and Textures [9], which are same as Sun et al. [65]) to fully evaluate the method’s performance across various semantic and scenario contexts. (2) We evaluated our method on various small-scale datasets. Specifically, we considered the following ID datasets: FashionMNIST [76], STL10 [10], OxfordIIIPet [55], Food-101 [6], CUB-200 [69], PlantVillage [30], LFW [27], Stanford-dogs [34], FGVC-Aircraft [47], Grocery Store [36], and CIFAR-10 [37]. (3) We assessed our method on hard OOD tasks [74, 50]. Following the standards of the MCM [48], we evaluated using subsets of ImageNet-1k, namely ImageNet-10 and ImageNet-20, which have similar classes (e.g., dog (ID) vs. wolf (OOD)). During the experiments, we ensured that each OOD dataset did not overlap with the ID dataset in terms of classes.
@@
In our experiment, all algorithms uniformly employ CLIP [59] as the pre-trained model, which is one of the most prevalent and publicly available visual-linguistic models. Specifically, we utilize CLIP-B/16 as the foundational evaluation model, consisting of a ViT-B/16 transformer [16] serving as the image encoder and a masked self-attention Transformer [68] as the text encoder. Additionally, unless otherwise specified, the temperature coefficient is uniformly set to 1 across all algorithms.
@@ 4.2 Main Results
The benchmarking of large-scale OOD datasets demonstrates the feasibility of the method for real-world applications and holds significant value. Typically, we employ ImageNet-1k as the ID dataset for Large-scale OOD detection. Table  3 presents the performance of our method in comparison with other approaches under this benchmark. Overall, our method surpasses other methods, achieving superior performance. When juxtaposed against the average performance of Zero-shot methods, CMA demonstrates enhancements of 2.22% in terms of AUROC and 2.68% in terms of FPR95. Similarly, when juxtaposed against the average performance of Require training methods, CMA also demonstrates improvements, achieving enhancements of 0.79% in AUROC and 3.13% in FPR95, thereby showcasing its exceptional performance.
@@
In contrast to benchmarks based on large-scale OOD datasets, small-scale OOD detection often features fewer distinct ID categories. Demonstrating robust performance across these categories is indicative of a method’s scalability. Table  3 illustrates the efficacy of our approach across various ID datasets. A notable outcome is that our method achieves impressive performance across these datasets, especially when no specific training was tailored to each dataset.
Hard OOD detection.
@@
Table  3 presents the outcomes of various zero-shot approaches in hard out-of-distribution (OOD) detection. For dataset configurations, we alternated between Imagenet-10 and Imagenet-20 (See Appendix A), which share similar semantic content, as the ID and OOD datasets, respectively. Alternatively, we employed the setup of using Imagenet-10 and Imagenet-20 as the ID datasets, with Imagenet-100 serving as the OOD dataset. Additionally, to account for spurious correlations, we utilized the Spurious OOD detection benchmark introduced by Ming et al. [48] and conducted comparative experiments on the Waterbirds dataset [50]. Across various experiments, CMA outperformed MSP and MCM in terms of AUROC and FPR95.
4.3 Ablation study
Whether to use softmax scaling.
@@ 4.3 Ablation study
Although MCM has demonstrated the significant role of softmax scaling in CLIP-based OOD detection, our method CMA underscores the even more crucial role of softmax scaling. Table  5 illustrates the disparities between the use and non-use of softmax scaling. A crucial factor is that without softmax scaling, the score formula in CMA would become $\mathcal{S}_{\text{CMA}}^{w/o}=\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{\hat{y}})/\tau$, where $\hat{y}=\mathop{\arg\max}\limits_{1\leq i\leq N}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$ represents the most matching concept index in ID categories. In comparison to the original formula, the absence of softmax scaling essentially excludes Agents from the final score calculation, resulting in a score of $\mathcal{S}_{\text{CMA}}^{w/o}=\max_{i\in[N]}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})/\tau$. This prevents the construction of vector triangle relationships, thus negating the impact of Agents and leading to a significant reduction in performance. This also validates the rationality and effectiveness of $\mathcal{S}_{\text{CMA}}$.
Comparative performance across models.

### v2

@@ 4 Experiments
MSP (Hendrycks and Gimpel 2016) 40.89 88.63 65.81 81.24 67.90 80.14 64.96 78.16 59.89 82.04
Energy (Liu et al. 2020) 21.59 95.99 34.28 93.15 36.64 91.82 51.18 88.09 35.92 92.26
ODIN (Liang, Li, and Srikant 2017) 30.22 94.65 54.04 87.17 55.06 85.54 51.67 87.85 47.75 88.80
ViM (Wang et al. 2022b) 32.19 93.16 54.01 87.19 60.67 83.75 53.94 87.18 50.20 87.82
KNN (Sun et al. 2022) 29.17 94.52 35.62 92.67 39.61 91.02 64.35 85.67 42.19 90.97
NPOS (Tao et al. 2022) 16.58 96.19 43.77 90.44 45.27 89.44 46.12 88.80 37.93 91.22
CoOp (Zhou et al. 2022) 43.38 91.26 38.53 91.95 46.68 89.09 50.64 87.83 44.81 90.03
LoCoOp (Miyai et al. 2024) 38.49 92.49 33.27 93.67 39.23 91.07 49.25 89.13 40.17 91.53
@@
FashionMNIST (Xiao, Rasul, and Vollgraf 2017) 0.00 100.00 0.00 100.00 0.00 100.00 13.61 93.60 3.40 98.40
STL10 (Coates, Ng, and Lee 2011) 0.00 100.00 2.56 99.01 0.00 100.00 0.00 100.00 0.64 99.75
OxfordIIIPet (Parkhi et al. 2012) 0.00 99.89 0.37 99.89 0.96 99.71 0.35 99.85 0.42 99.84
Food101 (Bossard, Guillaumin, and Van Gool 2014) 0.25 99.90 0.49 99.86 1.79 99.40 2.99 99.33 1.38 99.62
CUB-200 (Wah et al. 2011) 0.00 100.00 0.00 100.00 0.00 99.99 0.00 100.00 0.00 100.00
PlantVillage (Hughes, Salathé et al. 2015) 2.95 97.83 0.18 98.11 1.12 98.41 4.37 97.56 2.16 97.98
LFW (Huang et al. 2012) 2.89 99.14 2.24 99.52 8.04 98.25 18.88 95.34 8.01 98.06
Stanford-dogs (Khosla et al. 2011) 0.11 99.92 0.16 99.91 0.58 99.73 0.67 99.72 0.38 99.82
FGVC-Aircraft (Maji et al. 2013) 0.00 99.99 0.67 99.87 1.02 99.69 0.00 99.99 0.42 99.89
Grocery Store (Klasson, Zhang, and Kjellström 2019) 0.03 99.89 0.15 99.97 0.69 99.82 0.79 99.76 0.42 99.86
CIFAR10 (Krizhevsky, Hinton et al. 2009) 0.00 100.00 5.12 98.29 2.56 99.67 0.00 99.88 1.92 99.46
CIFAR100 (Krizhevsky, Hinton et al. 2009) 12.80 97.15 17.92 95.66 15.36 96.19 4.53 98.42 12.65 96.86
@@ 4.1 Setup
We conducted a comprehensive evaluation of the performance of our method across various dimensions and compared it with widely employed OOD detection algorithms. (1) We assessed our approach on the ImageNet-1k OOD benchmark. This benchmark utilizes the large-scale visual dataset ImageNet-1k (Deng et al. 2009) as the ID data and four OOD datasets (including subsets of iNaturalist (Van Horn et al. 2018), SUN (Xiao et al. 2010), Places(Zhou et al. 2017), and Textures (Cimpoi et al. 2014), which are same as Sun et al. (Sun et al. 2022)) to fully evaluate the method’s performance across various semantic and scenario contexts. (2) We evaluated our method on various small-scale datasets. Specifically, we considered the following ID datasets: FashionMNIST (Xiao, Rasul, and Vollgraf 2017), STL10 (Coates, Ng, and Lee 2011), OxfordIIIPet (Parkhi et al. 2012), Food-101 (Bossard, Guillaumin, and Van Gool 2014), CUB-200 (Wah et al. 2011), PlantVillage (Hughes, Salathé et al. 2015), LFW (Huang et al. 2012), Stanford-dogs (Khosla et al. 2011), FGVC-Aircraft (Maji et al. 2013), Grocery Store (Klasson, Zhang, and Kjellström 2019), and CIFAR-10 (Krizhevsky, Hinton et al. 2009). (3) We assessed our method on hard OOD tasks (Winkens et al. 2020; Ming, Yin, and Li 2022). Following the standards of the MCM (Ming et al. 2022a), we evaluated using subsets of ImageNet-1k, namely ImageNet-10 and ImageNet-20, which have similar classes (e.g., dog (ID) vs. wolf (OOD)). During the experiments, we ensured that each OOD dataset did not overlap with the ID dataset in terms of classes.
@@
In our experiment, all algorithms uniformly employ CLIP (Radford et al. 2021) as the pre-trained model, which is one of the most prevalent and publicly available visual-linguistic models. Specifically, we utilize CLIP-B/16 as the foundational evaluation model, consisting of a ViT-B/16 transformer (Dosovitskiy et al. 2020) serving as the image encoder and a masked self-attention Transformer (Vaswani et al. 2017) as the text encoder. Additionally, unless otherwise specified, the temperature coefficient is uniformly set to 1 across all algorithms.
@@ 4.2 Main Results
The benchmarking of large-scale OOD datasets demonstrates the feasibility of the method for real-world applications and holds significant value. Typically, we employ ImageNet-1k as the ID dataset for Large-scale OOD detection. Table  2 presents the performance of our method in comparison with other approaches under this benchmark. Overall, our method surpasses other methods, achieving superior performance. When juxtaposed against the average performance of Zero-shot methods, CMA demonstrates enhancements of 2.22% in terms of AUROC and 2.68% in terms of FPR95. Similarly, when juxtaposed against the average performance of Require training methods, CMA also demonstrates improvements, achieving enhancements of 0.79% in AUROC and 3.13% in FPR95, thereby showcasing its exceptional performance.
@@
In contrast to benchmarks based on large-scale OOD datasets, small-scale OOD detection often features fewer distinct ID categories. Demonstrating robust performance across these categories is indicative of a method’s scalability. Table  2 illustrates the efficacy of our approach across various ID datasets. A notable outcome is that our method achieves impressive performance across these datasets, especially when no specific training was tailored to each dataset.
5 Discussion
Figure 4: Impact curve of $k=\frac{\text{number of Agents}}{\text{number of ID Labels}}$ on the performance of CMA. Left shows that as $k$ gradually increases, the FPR95 on various datasets generally decreases, with the fastest decline occurring in the range of $k$ less than 0.5, followed by a gradual slowdown, which is more evident on the average curve. Right shows that as $k$ gradually increases, the AUROC gradually increases, also with a rapid rise followed by a gradual slowdown.
Figure 5: Comparison with different Agents. Left shows the performance of different agents on various datasets in terms of FPR95, while Right shows the performance in terms of AUROC. Clearly, the agent that performs best on average across all datasets does not perform best on every dataset. Moreover, even agents that perform poorly on average can still show decent performance on certain datasets.
The number of Agents.
In section 3, we present the score calculation formula $\mathcal{S}_{\text{CMA}}(\mathbf{x}^{\prime};\mathcal{Y}_{in},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I})$ for CMA, where we set the number of Agents to $N$, aligning with the number of ID categories. This number was determined based on extensive experiments, with $k$ set to $k=\frac{M}{N}$, where $M$ is the number of Agents.Concurrently, the formula for calculating the score has been modified to $\mathcal{S}_{\text{CMA}}^{\prime}(\mathbf{x}^{\prime};\mathcal{Y}_{in},\mathcal{Y}_{ntc},\mathcal{T},\mathcal{I})=\frac{\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{\hat{y}})/\tau)}{\sum_{i=1}^{N}\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})/\tau)+\sum_{j=1}^{M}\exp(\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{j+N})/\tau)}$, where $\hat{y}=\mathop{\arg\max}\limits_{1\leq i\leq N}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$. Experiments were conducted for various values of $k$. Figure 4 illustrates the results of our experiments conducted on ImageNet-1k. The optimal value of $k$ was identified as 1. Additionally, it was observed that an excessive number of Agents does not enhance the overall performance of CMA, whereas a insufficient number of Agents can adversely affect it. This finding illuminates the rationality of the vector triangle relationship. An insufficient number of Agents fails to exert the same impact on OOD data as it does on ID categories, leading to structural disarray and performance degradation. Once the number of Agents reaches a certain threshold, a stable structure is established, after which increasing the number of Agents does not significantly improve performance.
Performance of different Agents.
Figure 5 illustrates the impact of different neutral terms on the OOD detection performance of agents. It is evident that the impact of different agents on performance is substantial. On average, the best and worst outcomes in terms of FPR95 and AUROC can differ by 19.95% and 4.50%, respectively. From another perspective, certain agents excel in specific contexts, such as ”a large photo of the scene we live in.” While their overall performance across multiple environments lags significantly behind the optimal level, they achieve their best results when using the SUN dataset (Xiao et al. 2010) as OOD data. On one hand, this highlights a potential limitation of CMA, where different agents can yield disparate outcomes. On the other hand, it underscores the significant potential of CMA, as it can tailor specific agents to suit diverse environments, thereby enhancing performance.
6 Conclusion
Acknowledgements
This work was supported by National Natural Science Foundation of China, Grant Number: 62476109, 62206108, 62176184, and the Natural Science Foundation of Jilin Province, Grant Number: 20240101373JC, and Jilin Province Budgetary Capital Construction Fund Plan, Grant Number: 2024C008-5, and Research Project of Jilin Provincial Education Department, Grant Number: JJKH20241285KJ.
Appendix
Appendix A Cause Analysis and Statistical Analysis of Three Phenomena
A.1 Impact of prompt length on matching score
It is challenging to directly confirm the relationship between prompt length and matching score, as the matching score is obviously affected by multiple factors. A feasible method is to use only placeholders as prompts, such as ”xx xx”, which reduces the influence of the prompt itself and focuses on the length of the prompt. Therefore, we conducted a statistical experiment (Figure 6) using the CIFAR-10 as image inputs, with different lengths of the same placeholder as prompt inputs, and then averaged all the scores to obtain a statistical score distribution. The results clearly indicate a positive correlation between prompt length and matching score.
Figure 6: Distribution of scores for prompts of different lengths with different placeholders
Based on the above understanding, we define $S$ as a function of $L$, denoted as $S(L)$, and assume that within the range $R$, $S(L)$ is a monotonically increasing function. A simple linear model is proposed:
$S(L)=aL+b,\quad\text{for }L\in R$
where $a>0$ ensures that the function is increasing. To validate this relationship statistically, we employ a linear regression model using a dataset of $n$ prompts and their corresponding scores:
$S_{i}=\beta_{0}+\beta_{1}L_{i}+\varepsilon_{i},\quad i=1,2,\ldots,n$
where $\beta_{0}$ is the intercept, $\beta_{1}$ is the slope coefficient representing the effect of prompt length on the prediction score, and $\varepsilon_{i}$ is the error term. The significance of $\beta_{1}$ is tested using a t-test:
$t=\frac{\hat{\beta_{1}}}{SE(\hat{\beta_{1}})}$
where $\hat{\beta_{1}}$ is the estimated value of $\beta_{1}$ and $SE(\hat{\beta_{1}})$ is its standard error. If $|t|>t_{\text{crit}}$, where $t_{\text{crit}}$ is the critical value from the t-distribution at a chosen significance level, we reject the null hypothesis ($H_{0}:\beta_{1}=0$) and conclude that prompt length significantly affects the prediction score. Furthermore, a positive $\beta_{1}$ would support our hypothesis of a positive relationship between prompt length and prediction performance.
Figure 7: T-test results on the distribution of Absolute scores for prompts of different lengths
We conducted more extensive experiments on the larger dataset ImageNet-1k, selecting six placeholders ”x”, ”xx”, ”xxx”, ”thing”, ”time”, and ”the”. We removed softmax scaling and only performed more direct similarity matching. The matching results were called ”Absolute Score”, and then t-test was performed. The results (Figure 7) showed that ”Length significantly affects Score”.
A.2 Weight of words in text descriptions
We conducted statistical experiments on specific categories of CIFAR-10 using different prompts (Figure 8). For example, we used ”white cat”, ”black cat”, and ”a cat” as prompts, and images labeled as ”cat” in CIFAR-10 as image inputs to obtain a large number of matching scores, and finally averaged them to obtain a score distribution. It can be seen that although all three prompts represent ”cat”, and ”a something” should have stronger universality, there are still a large number of inputs that tend to be ”white” and ”black”, indicating the important influence of key words on matching scores.
Figure 8: Scores on different keywords
Let $P$ represent the set of all words in a given textual description, and $W_{i}$ denote the weight of the $i$-th word. We propose that the overall weight of the description, $W(P)$, can be expressed as a function of individual word weights and their associated attributes:
$W(P)=\sum_{i=1}^{n}W_{i}\cdot f_{i}(P)$
where $n$ is the total number of words in the description, and $f_{i}(P)$ is a multifaceted function associated with the $i$-th word. This function encapsulates various linguistic and contextual attributes, including but not limited to:
  • Word frequency: The occurrence rate of the word within the description and in a broader corpus.
  • Term frequency-inverse document frequency (TF-IDF): A numerical statistic reflecting the word’s importance in the document relative to a collection of documents.
  • Semantic significance: The word’s relevance to the overall meaning or topic of the description.
  • Syntactic role: The word’s grammatical function within the sentence structure.
  • Position: The location of the word within the description, potentially indicating its importance.
A.3 Impact of neutral prompts on images in ID and OOD inputs
Verifying the minimal impact of neutral prompt words on ID inputs is not difficult: We designed two sets of statistical experiments (Figure 9), one using CIFAR-10 as ID inputs, where the prompts were the labels of CIFAR-10, and the other using CIFAR-10 as OOD inputs, where the prompts were labels similar to those of CIFAR-10 in CIFAR-100. By gradually increasing the number of agents (neutral prompts), we observed the trend of score distribution changes. The results are clear: when faced with OOD inputs, the scores for each label decrease rapidly as the number of agents increases, while for ID inputs, the decrease is slower.
Figure 9: Score of different labels with different agent number
Formally, let us define the following sets:
- $ID$: the set of in-distribution images, i.e., images that belong to the categories the labels include
- $OOD$: the set of out-of-distribution images, i.e., images from categories not included in the labels
- $N$: the set of neutral prompts, defined as prompts that do not contain specific information about image content
We introduce $\Delta S(i,p)$ as the change in recognition score for image $i$ when prompt $p$ is applied. Our hypothesis posits two key relationships:
For in-distribution images:
$\forall i\in ID,\forall n\in N:|\Delta S(i,n)|\leq\epsilon$
where $\epsilon$ is a small positive constant representing the threshold for negligible change.
For out-of-distribution images:
$\forall i\in OOD,\forall n\in N:\Delta S(i,n)<-\delta$
where $\delta$ is a positive constant representing a significant decrease in score.
These relationships can be formalized as:
$P(|\Delta S(i,n)|\leq\epsilon|i\in ID,n\in N)\geq 1-\alpha$
$P(\Delta S(i,n)<-\delta|i\in OOD,n\in N)\geq 1-\beta$
where $\alpha$ and $\beta$ are small positive constants representing the acceptable error rates for the respective hypotheses. This formulation allows for statistical testing of the hypotheses using empirical data.
This can be verified by calculating the distribution of the differences between ID inputs and OOD inputs with and without an agent (Figure 10). Clearly, the distribution of differences for ID inputs is smaller and more concentrated, while that for OOD inputs is larger and more dispersed.
Figure 10: Comparison of difference distribution: ID inputs vs OOD inputs
Appendix B Method for Obtaining Neutral Prompts
The creation of neutral prompts through manual construction is feasible to a certain extent, by selecting words devoid of actual meaning to compose a short sentence. However, it is evident that employing large language models for the generation of neutral prompts is a superior option, as these models can generate a greater diversity of prompts, distinguishing themselves from manual constructions. We have developed two versions of prompts (See H.5 and H.6) to generate neutral prompts by large language models, the former being more stable, producing similar neutral prompts, while the latter is more diverse, capable of generating a wider array of prompts. The neutral prompts used and mentioned in this paper and in the experiments are all obtained through manual selection after generation by the large language models.
Appendix C Near OOD Detection Task
Table 3: Comparison results on near OOD detection tasks.
FPR95 / AUROC FPR95 / AUROC FPR95 / AUROC FPR95 / AUROC FPR95 / AUROC
MSP (Hendrycks and Gimpel 2016) 9.38 / 98.31 12.51 / 97.70 5.82 / 98.88 14.14 / 95.73 39.57 / 90.99
MCM (Ming et al. 2022a) 5.00 / 98.71 12.91 / 98.09 3.70 / 99.09 13.16 / 96.32 5.87 / 98.36
CMA(Ours) 3.1 / 99.19 6.2 / 98.71 2.94 / 99.18 10.32 / 96.64 3.22 / 99.01
@@ Appendix C Near OOD Detection Task
Table  3 presents the outcomes of various zero-shot approaches in near out-of-distribution (OOD) detection. For dataset configurations, we alternated between Imagenet-10 and Imagenet-20 (See Appendix H), which share similar semantic content, as the ID and OOD datasets, respectively. Alternatively, we employed the setup of using Imagenet-10 and Imagenet-20 as the ID datasets, with Imagenet-100 serving as the OOD dataset. Additionally, to account for spurious correlations, we utilized the Spurious OOD detection benchmark introduced by Ming et al. (Ming et al. 2022a) and conducted comparative experiments on the Waterbirds dataset (Ming, Yin, and Li 2022). Across various experiments, CMA outperformed MSP and MCM in terms of AUROC and FPR95.
Appendix D Ablation Study
D.1 Whether to use softmax scaling
@@
Although MCM has demonstrated the significant role of softmax scaling in CLIP-based OOD detection, our method CMA underscores the even more crucial role of softmax scaling. Table  5 illustrates the disparities between the use and non-use of softmax scaling. A crucial factor is that without softmax scaling, the score formula in CMA would become:
$\mathcal{S}_{\text{CMA}}^{w/o}=\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{\hat{y}})/\tau,$
@@
where $\hat{y}=\mathop{\arg\max}\limits_{1\leq i\leq N}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})$ represents the most matching concept index in ID categories. In comparison to the original formula, the absence of softmax scaling essentially excludes Agents from the final score calculation, resulting in a score of $\mathcal{S}_{\text{CMA}}^{w/o}=\max_{i\in[N]}\text{sim}(\mathbf{v}^{\prime},\mathbf{c}_{i})/\tau$. This prevents the construction of vector triangle relationships, thus negating the impact of Agents and leading to a significant reduction in performance. This also validates the rationality and effectiveness of $\mathcal{S}_{\text{CMA}}$.
D.2 Comparative performance across models